# Revisiting Applicable and Comprehensive Knowledge Tracing in Large-Scale Data

### v1

@@ A.3.2 Matrix Long Short-Term Memory
Similarly, to stabilize the exponential function in mLSTM, mLSTM employs the same stabilization technique as sLSTM (see Eq. 12 in Appendix A.3). The design of mLSTM supports highly parallelized processing, which not only improves computational efficiency but also allows the model to scale better to large datasets.
In addition, xLSTM introduces residual networks He et al. (2016) to stack sLSTM or mLSTM, enabling xLSTM to effectively process complex sequential data while improving the training stability of the model in deep networks.
A.4 Dataset Description and Processing Methods
@@
We conduct extensive experiments on three of the latest large-scale benchmark datasets from different platforms: (i) Assist17555https://sites.google.com/view/assistmentsdatamining/dataset?authuser=0 is the latest subset of the ASSISTments dataset released by Worcester Polytechnic Institute. ASSISTments is an online tutoring system that provides mathematics instruction and access services for students, widely used in mathematics courses for 4th to 12th-grade students in the United States. A key feature of ASSISTments is providing students with immediate feedback, allowing them to know whether their answers are correct after responding to questions. (ii) EdNet [^6] https://github.com/riiid/ednet is a substantial educational dataset collected by Santa, a multi-platform artificial intelligence tutoring service. Collected over two years, this dataset encompasses a wide range of student-system interactions across Android, iOS, and web platforms in Korea. It contains over 130 million learning interactions from approximately 780,000 students, making it one of the largest publicly available interactive education system datasets. The dataset is notable for its scale and hierarchical structure, offering rich insights into student activities and learning patterns. To ensure computational efficiency, we randomly selected 20,000 students from EdNet, similar to previous studies Liu et al. (2020); Lee et al. (2022); Cui et al. (2023). (iii) Comp [^7] https://github.com/wahr0411/PTADisc, which is part of PTADisc, is specifically selected for KT tasks in computational thinking courses. PTADisc originates from PTA, an online programming teaching assistant system developed by PTA Educational Technology Co., Ltd. for universities and society, based on students. PTADisc is currently the largest dataset in the field of personalized learning, which also includes different courses of varying data scales, providing options for various types of learning.
Following the data preprocessing method in CL4KT Lee et al. (2022), we exclude students with fewer than five interactions and all interactions involving unnamed concepts. Since a single question may involve multiple concepts, we convert the unique concept combinations within a single question into a new concept. The statistics after processing are shown in Table 5.
@@
A.5 Baseline Description
Here is a detailed description of the 17 baselines from 8 different categories in our experiment.
@@ A.5 Baseline Description
    • DKT Piech et al. (2015): DKT is a pioneering model that utilizes Recurrent Neural Networks (RNNs), specifically a single-layer Long Short-Term Memory (LSTM) network, to directly model students’ learning processes and predict their performance.
    • DKT+ Yeung and Yeung (2018): DKT+ is an enhanced version of DKT. It addresses the reconstruction and prediction inconsistency issues present in the DKT by introducing additional regularization terms to the loss function.
    • DKT-F Nagatani et al. (2019): DKT-F improves upon DKT by incorporating students’ forgetting behaviors into the modeling process.
@@
    • SAKT Pandey and Karypis (2019): SAKT leverages self-attention networks to analyze and understand the complex relationships between concepts and a student’s historical interactions with learning materials.
    • AKT Ghosh et al. (2020): AKT is an advanced KT model that incorporates a Rasch model to regularize concept and question embeddings and a modified Transformer architecture with adaptive attention weights computed by a distance-aware exponential decay to account for the time distance between questions and students’ previous interactions.
    • simpleKT Liu et al. (2023b): simpleKT is a simple but tough-to-beat baseline to KT that combines simplicity with robust performance.
    • FoLiBiKT Im et al. (2023): FoLiBi enhances attention-based KT models by incorporating a forgetting-aware linear bias mechanism. We introduce FoLiBi with AKT, namely FoLiBiKT.
    • sparseKT Huang et al. (2023): sparseKT employs a k-selection module with soft-thresholding sparse attention (sparseKT-soft) and top-K sparse attention (sparseKT-topK) to focus on high-attention items, ensuring efficient and focused attention on the most relevant items.
    • DTransformer Yin et al. (2023): DTransformer integrates question-level mastery with knowledge-level diagnosis through the use of Temporal and Cumulative Attention (TCA) and multi-head attention mechanisms. Additionally, a contrastive learning-based algorithm is used for enhancing the stability of the knowledge state diagnosis process.
@@
    • Mamba4KT Cao and Zhang (2024): By leveraging Mamba, a state-space model supporting parallelized training and linear-time inference, Mamba4KT achieves efficient resource utilization, balancing time and space consumption.
@@
    • GKT Nakagawa et al. (2019): GKT revolutionizes the traditional KT task by employing Graph Neural Networks (GNNs) to represent the relationships between concepts as a graph.
@@
    • DKVMN Zhang et al. (2017): DKVMN employs a static key matrix to capture the interrelationships among latent concepts and a dynamic value matrix for continuously updating and predicting a student’s knowledge mastery in real-time.
    • SKVMN Abdelrahman and Wang (2019): SKVMN combines recurrent modeling of DKT with memory networks of DKVMN to enhance tracking of learners’ knowledge states over time.
@@
    • ATKT Guo et al. (2021): ATKT is an attention-based LSTM model that employs adversarial training techniques to enhance generalization and reduce overfitting by applying perturbations to student interaction sequences.
@@
    • CL4KT Lee et al. (2022): CL4KT employs contrastive learning on augmented learning histories to enhance representation learning by distinguishing between similar and dissimilar student learning patterns.
@@
    • Deep-IRT Yeung (2019): Deep-IRT is an explainable KT model that combines the DKVMN with Item Response Theory (IRT) to provide detailed insights into learner trajectories and concept difficulties, bridging deep learning capabilities with psychometric interpretability.
    • AT-DKT Liu et al. (2023a): AT-DKT enhances the original DKT by incorporating two auxiliary learning tasks: one focused on predicting question tags and the other on evaluating individualized prior knowledge.
Figure 8: Structural sketch of AKT.
Datasets Settings Metrics AKT simpleKT FoLiBiKT sparseKT DTransformer DKVMN CL4KT Deep-IRT
@@
A.6 Input and Output Settings
We provide a detailed explanation of the input and output settings for DLKT models, including weakly applicable input and comprehensive output settings, as well as strongly applicable input and comprehensive output settings.
A.6.1 Weakly Applicable Input and Comprehensive Output Settings
We use AKT as an example to describe the common weakly applicable input and comprehensive output settings in DLKT models. Figure 8 shows a structural sketch of AKT. Clearly, AKT takes both historical interactions and future interactions as input, ignoring historical information through attention masking while representing knowledge learned up to the current time step through offset (right-shifting values in attention), and directly predicts questions at each time step. From this, we can see that although AKT’s setup is reasonable and does not lead to future information leakage, this input setting, while convenient, also causes complications in engineering implementation (engineering often requires cumbersome representation of future information as a padding value, and this common processing method does not seem suitable for KT, as KT tasks typically involve predicting future questions $12\sim t+$ 1 based on historical interactions $1\sim t$). We provide a solution to this below.
A.6.2 Strongly Applicable Input and Comprehensive Output Settings
Similarly, using AKT as an example, we describe how it can be transformed into a strongly applicable input and comprehensive output setting. Like DKT2, as shown in Figure 2 in the main text, the transformed AKT only takes the historical interactions $\{(q_{i},c_{i},r_{i})\}_{i=1}^{t}$ as input, with everything else remaining unchanged (note that the right-shift operation still needs to be retained because the attention does not mask the knowledge of the current time step). Before outputting the predicted score $r_{i}^{c_{i}}$, it first concretizes knowledge into the knowledge of each concept (by converting the original dimensions into the number of concepts through an MLP) and then the comprehensive knowledge state is obtained through a Sigmoid function.
A.7 Additional Experimental Results
Dataset Assist17 EdNet Comp
@@ A.7 Additional Experimental Results
DKT 0.5841 0.5787 0.4913 0.6600 0.6225 0.4775 0.7091 0.8045 0.3806
SAKT 0.5596 0.5534 0.5048 0.6546 0.6198 0.4788 0.6994 0.8037 0.3824
AKT 0.6185 0.6040 0.4862 0.6649 0.6241 0.4765 0.7054 0.8039 0.3815
DKVMN 0.5730 0.5701 0.4964 0.6572 0.6205 0.4781 0.7050 0.8034 0.3817
ATKT 0.6205 0.6077 0.4836 0.6639 0.6229 0.4768 0.7111 0.8049 0.3802
CL4KT 0.5892 0.5911 0.4890 - - - 0.7044 0.8032 0.3820
DKT2 0.6174 0.6041 0.4872 0.6646 0.6243 0.4768 0.7064 0.8047 0.3810
@@
A.7.1 Multi-step Prediction Results
@@
A.7.2 Prediction Results with Varying History Lengths
Figure 6 and Figure 7 show the prediction performance of DKT2 and several representative baselines with different history lengths on the EdNet and Comp, respectively.
A.7.3 Prediction Results Under Three Input Settings
@@
A.7.4 Discussion on Multi-concept Prediction
A.7.5 Ablation Study on ACC and RMSE
Figure 9 shows the ablation study of DKT2 on ACC and RMSE.

### v2

@@ 0.A.3.2 Matrix Long Short-Term Memory
Similarly, to stabilize the exponential function in mLSTM, mLSTM employs the same stabilization technique as sLSTM (see Eq. 12). The design of mLSTM supports highly parallelized processing, which not only improves computational efficiency but also allows the model to scale better to large datasets.
In addition, xLSTM introduces residual networks [32] to stack sLSTM or mLSTM, enabling xLSTM to effectively process complex sequential data while improving the training stability of the model in deep networks.
0.A.4 Dataset Description and Processing Methods
@@
We conduct extensive experiments on three of the latest large-scale benchmark datasets from different platforms: (i) Assist17444https://sites.google.com/view/assistmentsdatamining/dataset?authuser=0 is the latest subset of the ASSISTments dataset released by Worcester Polytechnic Institute. ASSISTments is an online tutoring system that provides mathematics instruction and access services for students, widely used in mathematics courses for 4th to 12th-grade students in the United States. A key feature of ASSISTments is providing students with immediate feedback, allowing them to know whether their answers are correct after responding to questions. (ii) EdNet [^5] https://github.com/riiid/ednet is a substantial educational dataset collected by Santa, a multi-platform artificial intelligence tutoring service. Collected over two years, this dataset encompasses a wide range of student-system interactions across Android, iOS, and web platforms in Korea. It contains over 130 million learning interactions from approximately 780,000 students, making it one of the largest publicly available interactive education system datasets. The dataset is notable for its scale and hierarchical structure, offering rich insights into student activities and learning patterns. To ensure computational efficiency, we randomly selected 20,000 students from EdNet, similar to previous studies [52, 46, 17]. (iii) Comp [^6] https://github.com/wahr0411/PTADisc, which is part of PTADisc, is specifically selected for KT tasks in computational thinking courses. PTADisc originates from PTA, an online programming teaching assistant system developed by PTA Educational Technology Co., Ltd. for universities and society, based on students. PTADisc is currently the largest dataset in the field of personalized learning, which also includes different courses of varying data scales, providing options for various types of learning.
Following the data preprocessing method in CL4KT [46], we exclude students with fewer than five interactions and all interactions involving unnamed concepts. Since a single question may involve multiple concepts, we convert the unique concept combinations within a single question into a new concept. The statistics after processing are shown in Table 6.
@@
0.A.5 Baseline Description
Here is a detailed description of the 18 baselines from 8 different categories in our experiment.
@@ 0.A.5 Baseline Description
    • DKT [69]: DKT is a pioneering model that utilizes Recurrent Neural Networks (RNNs), specifically a single-layer Long Short-Term Memory (LSTM) network, to directly model students’ learning processes and predict their performance.
    • DKT+ [94]: DKT+ is an enhanced version of DKT. It addresses the reconstruction and prediction inconsistency issues present in the DKT by introducing additional regularization terms to the loss function.
    • DKT-F [64]: DKT-F improves upon DKT by incorporating students’ forgetting behaviors into the modeling process.
@@
    • SAKT [67]: SAKT leverages self-attention networks to analyze and understand the complex relationships between concepts and a student’s historical interactions with learning materials.
    • AKT [26]: AKT is an advanced KT model that incorporates a Rasch model to regularize concept and question embeddings and a modified Transformer architecture with adaptive attention weights computed by a distance-aware exponential decay to account for the time distance between questions and students’ previous interactions.
    • simpleKT [54]: simpleKT is a simple but tough-to-beat baseline to KT that combines simplicity with robust performance.
    • FoLiBiKT [42]: FoLiBi enhances attention-based KT models by incorporating a forgetting-aware linear bias mechanism. We introduce FoLiBi with AKT, namely FoLiBiKT.
    • sparseKT [40]: sparseKT employs a k-selection module with soft-thresholding sparse attention (sparseKT-soft) and top-K sparse attention (sparseKT-topK) to focus on high-attention items, ensuring efficient and focused attention on the most relevant items.
    • DTransformer [95]: DTransformer integrates question-level mastery with knowledge-level diagnosis through the use of Temporal and Cumulative Attention (TCA) and multi-head attention mechanisms. Additionally, a contrastive learning-based algorithm is used for enhancing the stability of the knowledge state diagnosis process.
    • stableKT [47]: stableKT excels in length generalization, delivering stable and consistent performance across both short and long student interaction sequences. It employs a multi-head aggregation module that integrates dot-product and hyperbolic attention to capture hierarchical relationships between questions and their associated concepts.
@@
    • Mamba4KT [10]: By leveraging Mamba, a state-space model supporting parallelized training and linear-time inference, Mamba4KT achieves efficient resource utilization, balancing time and space consumption.
@@
    • GKT [65]: GKT revolutionizes the traditional KT task by employing Graph Neural Networks (GNNs) to represent the relationships between concepts as a graph.
@@
    • DKVMN [96]: DKVMN employs a static key matrix to capture the interrelationships among latent concepts and a dynamic value matrix for continuously updating and predicting a student’s knowledge mastery in real-time.
    • SKVMN [1]: SKVMN combines recurrent modeling of DKT with memory networks of DKVMN to enhance tracking of learners’ knowledge states over time.
@@
    • ATKT [29]: ATKT is an attention-based LSTM model that employs adversarial training techniques to enhance generalization and reduce overfitting by applying perturbations to student interaction sequences.
@@
    • CL4KT [46]: CL4KT employs contrastive learning on augmented learning histories to enhance representation learning by distinguishing between similar and dissimilar student learning patterns.
@@
    • Deep-IRT [93]: Deep-IRT is an explainable KT model that combines the DKVMN with Item Response Theory (IRT) to provide detailed insights into learner trajectories and concept difficulties, bridging deep learning capabilities with psychometric interpretability.
    • AT-DKT [53]: AT-DKT enhances the original DKT by incorporating two auxiliary learning tasks: one focused on predicting question tags and the other on evaluating individualized prior knowledge.
Datasets Settings Metrics AKT simpleKT FoLiBiKT sparseKT DTransformer stableKT DKVMN CL4KT Deep-IRT
@@
[…]
0.A.6 Additional Experimental Results
@@
0.A.6.1 Multi-step Prediction Results
@@
0.A.6.2 Prediction Results with Varying History Lengths
Fig. 7 and Fig. 8 show the prediction performance of DKT2 and several representative baselines with different history lengths on the EdNet and Comp, respectively.
0.A.6.3 Prediction Results Under Three Input Settings
@@
0.A.6.4 Ablation Study on ACC and RMSE
Fig. 9 shows the ablation study of DKT2 on ACC and RMSE.